OpenStudy (jade27p):

Rationalize the denominator of fraction with numerator square root of negative 36 and denominator the quantity of 2 minus 3i plus the quantity 3 plus 2i. @phi

OpenStudy (phi):

\[ \frac{\sqrt{-36}}{2-3i+3+2i} \] ?

OpenStudy (jade27p):

yes and 2-3i is in parenthesis, so is 3+2i

OpenStudy (phi):

first, I assume you can simplify sqrt(-36) ?

OpenStudy (jade27p):

yes

OpenStudy (phi):

what is it ?

OpenStudy (jade27p):

would it be 6i/(2-3i)+(3+2i)

OpenStudy (phi):

yes, next in the bottom add real to real and imaginary to imaginary what do you get for the bottom ?

OpenStudy (jade27p):

5-i

OpenStudy (phi):

so now we have \[ \frac{6i}{5-i}\] multiply top and bottom by the complex conjugate of 5-i do you know how to do that ?

OpenStudy (jade27p):

idk

OpenStudy (phi):

the complex conjugate of 5-i is 5+i (you "switch" the sign of the imaginary part) we now do (5-i)(5+i) notice this is similar to (a+b)(a-b) = a^2 - b^2 (we can use FOIL to show this is what you get) can you do (5-i)(5+i) ?

OpenStudy (jade27p):

6+30i/25

OpenStudy (phi):

can you do (5-i)(5+i) ?

OpenStudy (jade27p):

26

OpenStudy (phi):

yes it is 5^2 - i^2 or 25 - -1 = 25+1 = 26 if you do a lot of these you remember to just square both numbers and add: 25+1 so you have \[ \frac{6i(5+i)}{26} \] next "distributre" the 6i (multiply each term inside by 6i)

OpenStudy (jade27p):

-6+30i

OpenStudy (phi):

so you have \[ \frac{-6+30i}{26} \] it looks like we can divided top and bottom by 2 and simplify to get \[ \frac{-3+15i}{13} \] or \[ -\frac{3}{13} + \frac{15}{13} i \]

OpenStudy (jade27p):

thank you